# ConDiSim: Conditional Diffusion Models for Simulation-Based Inference

### v1

@@ Abstract
We present a conditional diffusion model - ConDiSim, for simulation-based inference of complex systems with intractable likelihoods. ConDiSim leverages denoising diffusion probabilistic models to approximate posterior distributions, consisting of a forward process that adds Gaussian noise to parameters, and a reverse process learning to denoise, conditioned on observed data. This approach effectively captures complex dependencies and multi-modalities within posteriors. ConDiSim is evaluated across ten benchmark problems and two real-world test problems, where it demonstrates effective posterior approximation accuracy while maintaining computational efficiency and stability in model training. ConDiSim offers a robust and extensible framework for simulation-based inference, particularly suitable for parameter inference workflows requiring fast inference methods.
@@ 1 Introduction
Statistical inference of model parameters from empirical observations is a fundamental challenge in scientific research, enabling researchers to derive meaningful insights from complex simulation models. These parameters govern the behavior of simulators that replicate real-world phenomena, providing a bridge between theoretical constructs and empirical observations [Lavin et al., 2021]. Calibrating these parameters to ensure that simulator outputs align with observed data constitutes an inverse problem, formally defined within the framework of simulation-based inference (SBI) [Cranmer et al., 2020]. Solving this inverse problem involves addressing uncertainties arising from model stochasticity and potential multi-valuedness, where different sets of parameter values can produce similar observations or similar parameters may lead to varied outputs. Additionally, parameter inference becomes increasingly complex when simulators operate as ‘black boxes’ with intractable likelihood functions, rendering traditional likelihood-based Bayesian methods impractical [Sisson et al., 2018].
Recent advances in machine learning have introduced powerful generative models capable of approximating complex data distributions without explicit likelihood computations. Among these, diffusion models have emerged as a particularly promising class [Sohl-Dickstein et al., 2015]. Originally developed for image and signal processing, diffusion models iteratively denoise data starting from pure noise, effectively learning data distributions through a sequence of transformations. Compared to alternative generative frameworks like Variational Autoencoders (VAEs) [Kingma and Welling, 2014], Generative Adversarial Networks (GANs) [Goodfellow et al., 2014], and flow-based models [Rezende and Mohamed, 2015], diffusion models offer significant advantages. They avoid issues like posterior collapse in VAEs and mode collapse in GANs, and are not constrained by invertible transformations as in flow-based models, enabling them to better model complex, high-dimensional, and multi-modal distributions. Diffusion models also demonstrate superior training stability and come with theoretical convergence guarantees due to their foundation in stochastic differential equations [Ho et al., 2020]. Their flexible architectures and sampling processes allow practitioners to balance sample quality and generation speed, enhancing their applicability in various contexts, including SBI.
Applying diffusion models to SBI leverages their ability to implicitly learn the conditional distribution of simulator parameters given observations, circumventing the need for explicit likelihood evaluations. This is particularly advantageous for inverse problems involving black-box simulators. Moreover, diffusion models naturally support amortized inference: a single, globally trained model approximates the posterior distribution across multiple observations, eliminating the need for separate optimization for each new observation [Kingma and Welling, 2014]. This capability is especially valuable when rapid inference is required. However, integrating diffusion models into SBI presents challenges. Their iterative sampling process can be computationally intensive, and their performance is sensitive to hyperparameters like noise schedules and network architecture. Addressing these issues is crucial for enhancing the practical utility of diffusion-based SBI methods [Nichol and Dhariwal, 2021].
Contribution: We present a conditional diffusion model (ConDiSim) for simulation-based inference that balances model expressiveness with training efficiency and simplicity. Experiments on benchmark tasks and real-world datasets show the effectiveness of the approach in recovering complex posterior distributions, while incurring minimal computational overhead. Further, simulation-based calibration checks validate the reliability and robustness of the proposed parameter inference approach.
@@
The paper is organized as follows. Section 2 outlines the related work. In Section 3, we formally define the SBI problem, followed by a detailed explanation of the ConDiSim model in Section 3.1 and Section 3.2 details our amortized inference strategy. A comprehensive comparison on benchmark, and real-world test problems is presented in Section 4, with additional results provided in the Appendix. Section 5 discusses the results, while Section 6 summarizes the key findings and suggests future directions.
@@ 2 Related Work
Simulation-based inference (SBI) has advanced significantly in recent years, evolving from traditional likelihood-free methods to modern deep learning approaches. Classical techniques like Approximate Bayesian Computation (ABC) [Beaumont et al., 2002] and its sequential variants such as Sequential Monte Carlo-ABC (SMC-ABC) [Toni et al., 2009], provide mathematical convergence guarantees but often struggle with high-dimensional problems [Sisson et al., 2018]. Although neural network-learned summary statistics [Jiang et al., 2017, Wrede et al., 2022] have contributed to mitigating this limitation, they introduce additional computational complexity.
@@
Early advancements in neural network-based amortized inference for SBI were pioneered by Neural Posterior Estimation (NPE) [Papamakarios and Murray, 2016], which introduced conditional normalizing flows to model the posterior distribution, significantly improving accuracy and scalability over traditional methods. Sequential extensions like SNPE-A [Papamakarios and Murray, 2018], SNPE-B [Lueckmann et al., 2017], and SNPE-C (also known as Automatic Posterior Transformation or APT) [Greenberg et al., 2019] enhanced efficiency by iteratively refining the model to focus on relevant parameter regions. However, these sequential methods require multiple training rounds and are constrained by the invertible transformations required in flow-based models.
Alternative sequential algorithms have also been proposed, such as Sequential Neural Ratio Estimation (SNRE) [Hermans et al., 2020], which learns the likelihood-to-evidence ratio, and Sequential Neural Likelihood Estimation (SNLE) [Papamakarios et al., 2019], which approximates the likelihood function directly. Generative Adversarial Training for SBI (GATSBI) [Ramesh et al., 2022] leverages adversarial networks to model complex distributions without explicit likelihood calculations. While effective for challenging inference problems, GATSBI faces issues common to adversarial training, such as instability and mode collapse [Arjovsky and Bottou, 2017, Arjovsky et al., 2017], which can limit robustness and complicate training, and it struggles in low-dimensional spaces as noted in [Ramesh et al., 2022].
Recently, Simformer [Gloeckler et al., 2024] combined score-based diffusion models with transformer architectures to accurately learn complex, high-dimensional posteriors, offering high flexibility albeit with substantial computational overhead. In contrast, our proposed method, ConDiSim, utilizes denoising-based diffusion models with a simpler architecture, maintaining competitive performance while being computationally efficient. This makes it suitable for applications such as model exploration, that may involve numerous parameter inference problems with inference time being a key consideration [Wrede and Hellander, 2019].
@@
In this study, we evaluate CondiSim against four distinct simulation-based inference (SBI) approaches: GATSBI [Ramesh et al., 2022], APT [Greenberg et al., 2019], NPE [Papamakarios and Murray, 2016], and Simformer [Gloeckler et al., 2024]. These methods represent a diverse spectrum of generative modeling techniques—namely, flow-based, adversarial, diffusion, one-shot, and sequential frameworks—adapted for SBI. This comprehensive selection facilitates an evaluation that emphasizes efficiency and scalability rather than focusing solely on the accuracy of posterior approximations. For further details on deep learning methodologies applied to SBI, readers are referred to [Zammit-Mangion et al., 2024, Radev et al., 2023a].
@@ 3 Simulation-Based Inference
where $p(\mathbf{y}\mid\bm{\theta})$ is the likelihood function and $p(\bm{\theta})$ is the prior distribution. In many complex models, especially those involving stochastic simulators, the likelihood function $p(\mathbf{y}\mid\bm{\theta})$ is intractable or computationally expensive to evaluate, rendering traditional inference methods like Markov Chain Monte Carlo (MCMC) algorithms [Andrieu et al., 2003] infeasible due to their reliance on explicit likelihood evaluations.
When the likelihood function is inaccessible, a popular strategy is to generate empirical samples using a stochastic simulator $\mathcal{M}(\bm{\theta},\bm{\xi})$ that produces synthetic data $\mathbf{y}$ given parameters $\bm{\theta}$ and random variables $\bm{\xi}$ representing system noise. Sampling $\bm{\theta}$ from a prior distribution $p(\bm{\theta})$ informed by domain knowledge or chosen to be noninformative and generating the corresponding $\mathbf{y}$ through the simulator, we obtain samples from the joint distribution $p(\bm{\theta},\mathbf{y})$. These joint samples allow us to approximate the posterior distribution $p(\bm{\theta}\mid\mathbf{y})$ using generative modeling techniques that learn the inverse mapping from observations $\mathbf{y}$ to parameters $\bm{\theta}$ without requiring explicit likelihood calculations.
3.1 ConDiSim: A Conditional Diffusion Framework for Simulation-Based Inference
We now introduce ConDiSim — a denoising conditional diffusion model [Ho et al., 2020] that systematically addresses the challenge of intractable likelihoods by leveraging samples from the joint distribution $p(\bm{\theta},\mathbf{y})$ , enabling efficient and accurate posterior inference without the need for explicit likelihood evaluations. The model consists of a forward diffusion process that gradually adds Gaussian noise to the parameters $\bm{\theta}$, transforming them into a tractable noise distribution. The reverse process then denoises these parameters, conditioned on observed data $\mathbf{y}$, to approximate the posterior distribution $p(\bm{\theta}\mid\mathbf{y})$. Figure 1 provides a schematic overview of the ConDiSim model.
In the forward diffusion process, we construct a discrete-time Markov chain $\{\bm{\theta}_{t}\}_{t=0}^{T}$, initialized at $\bm{\theta}_{0}=\bm{\theta}$. At each step $t$, the state $\bm{\theta}_{t}$ is derived from $\bm{\theta}_{t-1}$ by perturbing it with Gaussian noise according to the transition kernel:
@@ 3.1 ConDiSim: A Conditional Diffusion Framework for Simulation-Based Inference
where $\beta_{t}\in(0,1)$ determines the noise level at each step, and $\mathbf{I}$ is the identity matrix. This step-wise transformation progressively corrupts $\bm{\theta}_{0}$, moving its mean towards zero and increasing its variance to unity, such that as $t\to T$, the state converges to $q(\bm{\theta}_{T})=\mathcal{N}(\mathbf{0},\mathbf{I})$. Alternatively, $\bm{\theta}_{t}$ can be directly computed from $\bm{\theta}_{0}$ using the marginal distribution:
@@
$p(\bm{\theta}_{0:T}\mid\mathbf{y})=p(\bm{\theta}_{T})\,\prod_{t=1}^{T}p(\bm{\theta}_{t-1}\mid\bm{\theta}_{t},\mathbf{y}),$ (5)
Training the reverse model involves maximizing the log-posterior of the latent variables given the observations:
@@
[…]
However, evaluating $\log p(\bm{\theta}_{0}\mid\mathbf{y})$ directly is intractable, as it requires integrating over the full trajectory of intermediate latent states. To address this, we employ a variational approach by introducing the joint forward distribution $q(\bm{\theta}_{1:T}\mid\bm{\theta}_{0})$. Using Jensen’s inequality, the lower bound on the log-posterior is given by:
@@
$\mathcal{L}_{\mathrm{ConDiSim}}=\mathcal{L}_{\mathrm{recon}}+\mathcal{L}_{\mathrm{prior}}+\mathcal{L}_{\mathrm{denoise}}$ (9)
where, the reconstruction term, $\mathcal{L}_{\mathrm{recon}}=\mathbb{E}_{q}[\log p(\bm{\theta}_{0}\mid\bm{\theta}_{1},\mathbf{y})]$, evaluates how well the original parameters are recovered from the final denoised state. The prior term, $\mathcal{L}_{\mathrm{prior}}=-D_{\mathrm{KL}}(q(\bm{\theta}_{T}\mid\bm{\theta}_{0})\,\|\,p(\bm{\theta}_{T}))$, enforces consistency with the prior at the final noised step. The denoising term, $\mathcal{L}_{\mathrm{denoise}}=-\sum_{t=2}^{T}\mathbb{E}_{q}[D_{\mathrm{KL}}(q(\bm{\theta}_{t-1}\mid\bm{\theta}_{t},\bm{\theta}_{0})\,\|\,p(\bm{\theta}_{t-1}\mid\bm{\theta}_{t},\mathbf{y}))]$, aligns reverse transitions with the forward process. For sufficiently large $T$, the reconstruction and prior terms vanish, leaving the denoising KL as the dominant training objective:
@@
To eliminate the explicit dependence on $\bm{\theta}_{0}$, we substitute $\bm{\theta}_{0}=(\bm{\theta}_{t}-\sqrt{1-\bar{\alpha}_{t}}\,\bm{\epsilon})/\sqrt{\bar{\alpha}_{t}}$ (with $\bm{\epsilon}\sim\mathcal{N}(\mathbf{0},\mathbf{I})$) into Eq. (3), thereby reparameterizing $\bm{\mu}_{q}$ to yield a noise-parameterized expression for the mean:
@@
Assuming the learned reverse process preserves the covariance structure ($\bm{\Sigma}_{p}=\bm{\Sigma}_{q}$), the KL divergence in Eq. (10) reduces to the squared Euclidean distance between the means:
@@
where, $\bm{\epsilon}_{\phi}(\bm{\theta}_{t},\mathbf{y},t)$ is neural network that, given the noisy latent $\bm{\theta}_{t}$, conditioning $\mathbf{y}$, and time step $t$, predicts the noise. Substituting the mean parameterizations from Eqs.(12) and (14) into the loss of Eq.(13) and discarding all constant factors, yields the compact noise–prediction objective:
$\mathcal{L}_{\text{ConDiSim}}=\mathbb{E}_{\,t,\,\bm{\epsilon},\,(\bm{\theta}_{0},\mathbf{y})\sim p(\theta,y)}\Bigl{[}\big{\lVert}\bm{\epsilon}-\bm{\epsilon}_{\phi}(\bm{\theta}_{t},\mathbf{y},t)\big{\rVert}^{2}\Bigr{]}.$ (15)
Figure 1: The ConDiSim Model: (a) The forward diffusion process incrementally adds Gaussian noise to the parameter vector ($\bm{\theta}$), generating a sequence of latent states ($\{\bm{\theta}_{t}\}$); (b) The reverse diffusion process reconstructs the posterior distribution ($p(\bm{\theta}\mid\mathbf{y})$) by leveraging the observed data ($\mathbf{y}$) for denoising.
3.2 Amortized Inference with ConDiSim
In our framework, we implement the reverse diffusion process as an instance of amortized inference. For a given observation $\mathbf{y_{o}}$, the reverse process is initialized by drawing $N$ latent variables $\bm{\theta}_{T}^{(i)}$ from the standard normal distribution $\mathcal{N}(\mathbf{0},\mathbf{I})$ for $i=1,\dots,N$. These latent samples are then conditioned on $\mathbf{y_{o}}$ and iteratively refined through the learned reverse diffusion process. Specifically, at each reverse time step, the transition is characterized by,
$\bm{\theta}_{t-1}^{(i)}\mid\bm{\theta}_{t}^{(i)},\bm{y_{o}}\sim\mathcal{N}\!\left(\bm{\mu}_{\phi}\!\left(\bm{\theta}_{t}^{(i)},\mathbf{y_{o}},t\right),\,\bm{\Sigma}_{t}\right),$
where $\bm{\mu}_{\phi}$ and $\bm{\Sigma}_{t}$ represent the learned mean and covariance parameters that determine the probabilistic mapping between consecutive latent states in the reverse process. After iterating from $t=T$ down to $t=0$, the process yields refined latent samples $\bm{\theta}_{0}^{(i)}$ that approximate the target posterior distribution $p(\bm{\theta}\mid\mathbf{y_{o}})$. This method efficiently produces high-quality posterior samples suitable for applications such as parameter estimation and uncertainty quantification.

### v2

@@ Abstract
We present ConDiSim, a conditional diffusion model for simulation-based inference in complex systems with intractable likelihoods. ConDiSim leverages denoising diffusion probabilistic models to approximate posterior distributions, consisting of a forward process that adds Gaussian noise to parameters, and a reverse process learning to denoise, conditioned on observed data. This approach effectively captures complex dependencies and multi-modalities within posteriors. ConDiSim is evaluated across ten benchmark problems and two real-world test problems, where it demonstrates effective posterior approximation accuracy while maintaining computational efficiency and stability in model training. ConDiSim provides a robust and extensible framework for simulation-based inference, well suited to parameter estimation tasks that demand fast methods for handling noisy, time series observations.
@@ 1 Introduction
Statistical inference of model parameters from empirical observations is a fundamental challenge in scientific research, enabling researchers to derive meaningful insights from complex simulation models. These parameters govern the behavior of simulators that replicate real-world phenomena, providing a bridge between theoretical constructs and empirical observations [Lavin et al., 2021]. Calibrating these parameters to ensure that simulator outputs align with observed data constitutes an inverse problem, formally defined within the framework of simulation-based inference (SBI) [Cranmer et al., 2020]. Such inverse problems are often ill-posed, meaning they may lack unique or stable solutions, especially when the simulator is stochastic or when different parameter configurations produce similar observations [Sisson et al., 2018]. To address these challenges, the Bayesian formulation of SBI places a prior distribution over the parameters and infers a posterior distribution conditioned on observed data. This posterior captures the full set of plausible solutions and enables principled uncertainty quantification, offering a more robust alternative to relying on single-point estimates.
Recent advances in machine learning have introduced powerful generative models capable of approximating complex data distributions without explicit likelihood computations. Among these, diffusion models have emerged as a particularly promising class [Sohl-Dickstein et al., 2015, Ho et al., 2020, Song et al., 2020]. They model data by learning to reverse a fixed noising process, gradually denoising samples from pure noise. Moreover, diffusion models naturally support amortized inference [Kingma and Welling, 2014]: a single trained model approximates posteriors for any observation without per-instance optimization, enabling faster inference for new data.
This work explores diffusion models for addressing common SBI challenges: limited simulation budgets from expensive simulators (particularly in systems biology and epidemiology with stiff ODE/PDE solvers), noisy and stochastic outputs, multimodal posteriors from nonlinearities and identifiability issues, and the need for architectures that integrate parameters, timesteps, and observations to recover diverse posterior structures. ConDiSim is demonstrated to be well-suited for SBI problems involving learning low-dimensional posteriors from noisy, high-dimensional discrete-time observations.
Contributions: (i)We adapt conditional Denoising Diffusion Probabilistic Models (DDPMs) for SBI to learn complex posterior distributions under limited simulation budgets; (ii) ConDiSim introduces a novel reverse diffusion architecture for SBI that employs time-dependent Feature-wise Linear Modulation (FiLM) to jointly condition on noisy parameters, diffusion timesteps, and observations for posterior estimation; (iii) We empirically evaluate ConDiSim systematically across benchmarks and real-world SBI tasks, analyzing performance under limited simulation budgets, noisy outputs, and multimodal posteriors.
@@
Outline: The paper is organized as follows. Section 2 outlines the related work. Section 3 formally defines the SBI problem. Section 3.1 introduces the ConDiSim model, and Section 3.2 details its conditioning mechanism and sampling procedure, including the proposed architecture. Section 4 presents a comprehensive evaluation on benchmark and real-world problems, with additional results provided in the Appendix. Section 5 presents a discussion of the results, while Section 6 concludes the paper.
@@ 2 Related Work
Simulation-based inference has advanced significantly in recent years, evolving from traditional statistical likelihood-free methods to contemporary deep learning frameworks. Classical techniques like Approximate Bayesian Computation (ABC) [Beaumont et al., 2002] and Sequential Monte Carlo-ABC (SMC-ABC) [Toni et al., 2009] offer theoretical convergence guarantees but may scale poorly to high dimensional problems [Sisson et al., 2018]. While neural summary statistics [Jiang et al., 2017, Wrede et al., 2022] can reduce dimensionality, they introduce trade-offs: additional computational burden and the possibility of losing information necessary for accurate posterior estimation.
@@
Neural Posterior Estimation (NPE) [Papamakarios and Murray, 2016] pioneered the use of deep generative models for amortized SBI, introducing conditional normalizing flows to model the posterior distribution, substantially improving accuracy and scalability over traditional methods. Sequential extensions like SNPE-A [Papamakarios and Murray, 2018], SNPE-B [Lueckmann et al., 2017], and SNPE-C (also known as Automatic Posterior Transformation or APT) [Greenberg et al., 2019] enhanced accuracy by iteratively refining the model to focus on relevant parameter regions. However, these sequential methods require multiple rounds of training and are constrained by the invertible transformations required in flow-based models. Recently, Flow Matching Posterior Estimation (FMPE) [Wildberger et al., 2023] introduced continuous normalizing flows based on conditional vector fields, overcoming architectural constraints and efficiently scaling to high-dimensional inference tasks. However, FMPE requires ODE solvers at inference time, introducing additional computational costs, and its performance depends critically on the choice of flow interpolant, which can be challenging to optimize under tight simulation budgets.
Beyond flow-based methods, other deep generative approaches to SBI include adversarial models. Generative Adversarial Training for SBI (GATSBI) [Ramesh et al., 2022] uses adversarial networks to model posteriors but suffers from training instability, mode collapse [Arjovsky and Bottou, 2017, Arjovsky et al., 2017], and poor performance in low-dimensional settings [Ramesh et al., 2022]. More recently, Simformer [Gloeckler et al., 2024] combined score-based diffusion models with a transformer architecture, enabling high-fidelity posterior estimation but at the expense of increased computational cost. Score-based diffusion models have also been explored in the tall-data setting of SBI, where multiple observations are combined efficiently using amortized score networks [Linhart et al., 2024]. In contrast, ConDiSim uses DDPMs with a lightweight, FiLM-based architecture that avoids heavy attention mechanisms while achieving competitive performance with significantly lower inference and training overhead. This makes it especially suitable for settings like iterative model exploration, where many inference queries must be resolved quickly [Wrede and Hellander, 2019].
@@
In this study, we evaluate ConDiSim against five distinct SBI approaches: NPE [Papamakarios and Murray, 2016], APT [Greenberg et al., 2019], GATSBI [Ramesh et al., 2022], FMPE [Wildberger et al., 2023], and Simformer [Gloeckler et al., 2024]. Together, these methods span a wide spectrum of generative modeling paradigms, including discrete flow-based methods in both one-shot and sequential forms, continuous flow-based approaches, adversarial models, and diffusion models in both discrete and continuous formulations. This selection enables a balanced evaluation that emphasizes efficiency and scalability in addition to posterior accuracy. For a broader overview of deep learning approaches to SBI, see [Zammit-Mangion et al., 2024, Radev et al., 2023a].
@@ 3 Simulation-Based Inference
where $p(\mathbf{y}\mid\bm{\theta})$ is the likelihood and $p(\bm{\theta})$ is the prior distribution. In many scientific models, particularly those involving stochastic simulators, the likelihood function $p(\mathbf{y}\mid\bm{\theta})$ is either intractable or computationally expensive to evaluate, rendering likelihood-based methods such as Markov Chain Monte Carlo (MCMC) infeasible [Andrieu et al., 2003].
When the likelihood is inaccessible, a common strategy is to sample parameters from a prior, $\bm{\theta}\sim p(\bm{\theta})$, where $p(\bm{\theta})$ may be informed by domain knowledge or chosen to be noninformative, such as a broad uniform prior. For each sampled parameter, a corresponding observation is generated via the simulator, $\mathbf{y}\sim p(\mathbf{y}\mid\bm{\theta})$. This yields joint samples $(\bm{\theta},\mathbf{y})\sim p(\bm{\theta},\mathbf{y})$, which can then be used to train generative models that approximate the posterior $p(\bm{\theta}\mid\mathbf{y})$ without requiring explicit likelihood evaluation.
3.1 ConDiSim: Conditional Diffusion Framework for SBI
We adapt conditional denoising diffusion probabilistic models [Ho et al., 2020] to SBI, where the goal is to approximate the posterior $p(\bm{\theta}\mid\mathbf{y})$ in likelihood-free settings using joint samples $(\bm{\theta},\mathbf{y})\sim p(\bm{\theta},\mathbf{y})$. The model employs a forward diffusion process that gradually corrupts the parameters $\bm{\theta}$ with Gaussian noise until they follow a tractable prior distribution. The reverse process then learns to denoise these corrupted parameters, conditioned on observations $\mathbf{y}$, thereby recovering samples from the posterior distribution $p(\bm{\theta}\mid\mathbf{y})$.
In the forward diffusion process, we construct a discrete-time Markov chain $\{\bm{\theta}_{t}\}_{t=0}^{T}$, initialized at $\bm{\theta}_{0}=\bm{\theta}$. At each step $t$, the new state $\bm{\theta}_{t}$ is derived from the old state $\bm{\theta}_{t-1}$ by perturbing it with Gaussian noise according to the transition kernel:
@@ 3.1 ConDiSim: Conditional Diffusion Framework for SBI
where $\beta_{t}\in(0,1)$ determines the noise level at each step, and $\mathbf{I}$ is the identity matrix. This step-wise transformation progressively corrupts $\bm{\theta}_{0}$, moving its mean towards zero and increasing its variance to unity, such that as $t\to T$, the state converges to $q(\bm{\theta}_{T})=\mathcal{N}(\mathbf{0},\mathbf{I})$.
Alternatively, $\bm{\theta}_{t}$ can be directly computed from $\bm{\theta}_{0}$ using the marginal distribution:
@@
$p(\bm{\theta}_{0:T}\mid\mathbf{y})=p(\bm{\theta}_{T})\,\prod_{t=1}^{T}p(\bm{\theta}_{t-1}\mid\bm{\theta}_{t},\mathbf{y}).$ (5)
Training the reverse model involves maximizing the log‑posterior of the latent variables, given the observations:
@@
$\text{where,}\quad\log p(\bm{\theta}_{0}\mid\mathbf{y})=\log\int p(\bm{\theta}_{0:T}\mid\mathbf{y})\,d\bm{\theta}_{1:T}.$ (7)
However, evaluating $\log p(\bm{\theta}_{0}\mid\mathbf{y})$ directly is intractable, as it entails integrating over the full trajectory of intermediate latent states. Instead, we employ a variational approach by introducing the joint forward distribution $q(\bm{\theta}_{1:T}\mid\bm{\theta}_{0})$. Using Jensen’s inequality, the lower bound on the log-posterior is:
@@
$\mathcal{L}_{\mathrm{ConDiSim}}=\mathcal{L}_{\mathrm{recon}}+\mathcal{L}_{\mathrm{prior}}+\mathcal{L}_{\mathrm{denoise}},$ (9)
where, the reconstruction term, $\mathcal{L}_{\mathrm{recon}}=\mathbb{E}_{q}[\log p(\bm{\theta}_{0}\mid\bm{\theta}_{1},\mathbf{y})]$, evaluates how well the original parameters are recovered from the final denoised state. The prior term, $\mathcal{L}_{\mathrm{prior}}=-D_{\mathrm{KL}}(q(\bm{\theta}_{T}\mid\bm{\theta}_{0})\,\|\,p(\bm{\theta}_{T}))$, enforces consistency with the prior at the final noised step. The denoising term, $\mathcal{L}_{\mathrm{denoise}}=-\sum_{t=2}^{T}\mathbb{E}_{q}[D_{\mathrm{KL}}(q(\bm{\theta}_{t-1}\mid\bm{\theta}_{t},\bm{\theta}_{0})\,\|\,p(\bm{\theta}_{t-1}\mid\bm{\theta}_{t},\mathbf{y}))]$, aligns reverse transitions with the forward process. For sufficiently large $T$, the reconstruction and prior terms vanish [Ho et al., 2020], leaving the denoising KL as the dominant training objective (i.e., $\mathcal{L}\approx\mathcal{L}_{\text{denoise}}$):
@@
To eliminate the explicit dependence on $\bm{\theta}_{0}$, we substitute $\bm{\theta}_{0}=(\bm{\theta}_{t}-\sqrt{1-\bar{\alpha}_{t}}\,\bm{\epsilon})/\sqrt{\bar{\alpha}_{t}}$ (with $\bm{\epsilon}\sim\mathcal{N}(\mathbf{0},\mathbf{I})$) into Eq. (3), reparameterizing $\bm{\mu}_{q}$ to yield a noise-parameterized expression for the mean:
@@
Assuming the learned reverse process preserves the covariance structure ($\bm{\Sigma}_{p}=\bm{\Sigma}_{q}$) [Ho et al., 2020], the KL divergence in Eq. (10) reduces to the squared Euclidean distance between the means:
@@
where, $\bm{\epsilon}_{\phi}(\bm{\theta}_{t},\mathbf{y},t)$ is a neural network that, given the noisy latent $\bm{\theta}_{t}$, conditioning $\mathbf{y}$, and time step $t$, predicts the noise. Substituting the mean parameterizations from Eqs.(12) and (14) into the loss of Eq.(13) and discarding all constant factors, yields the compact noise prediction objective:
$\mathcal{L}_{\text{ConDiSim}}\approx\mathbb{E}_{\,t,\,\bm{\epsilon},\,(\bm{\theta},\mathbf{y})\sim p(\bm{\theta},\mathbf{y})}\Bigl[\big\lVert\bm{\epsilon}-\bm{\epsilon}_{\phi}(\bm{\theta}_{t},\mathbf{y},t)\big\rVert^{2}\Bigr].$ (15)
Figure 1: ConDiSim Architecture: Each diffusion block is FiLM-modulated by timestep and observation embeddings to estimate the noise injected at step $t$ during forward diffusion.
3.2 Conditioning and Sampling
The objective in Eq. (15) requires a denoising network that predicts the noise $\bm{\epsilon}$ from the noisy state $\bm{\theta}_{t}$, conditioned on the observation $\mathbf{y}$ and the diffusion timestep $t$. Preserving this dual conditioning throughout the denoising trajectory is essential to align with $p(\bm{\theta}\mid\mathbf{y})$. A naive scheme that feeds $\bm{\theta}_{t};\,\mathbf{y};\,t]$ into a standard network often performs poorly in SBI, where $\dim(\mathbf{y})\!\gg\!\dim(\bm{\theta})$, input-level concatenation forces the network to disentangle $\mathbf{y}$ and $t$ from $\bm{\theta}_{t}$ in a single step, yielding an ill-conditioned coupling that dilutes the posterior signal and destabilizes training. To address this, the proposed ConDiSim (Fig. 1) model takes $(\bm{\theta}_{t},t,\mathbf{y})$ as inputs; while a condition encoder $f_{y}(\mathbf{y})$ and a timestep encoder $f_{t}(t)$ produce embeddings that are concatenated into a context vector $\mathbf{c}=[\,f_{y}(\mathbf{y});\;f_{t}(t)\,]$. A lightweight MLP maps $\mathbf{c}$ to per-block FiLM parameters $(\bm{\gamma}_{\ell}(\mathbf{c}),\bm{\beta}_{\ell}(\mathbf{c}))$, which modulate each diffusion block via $\operatorname{FiLM}_{\ell}(\mathbf{h}_{\ell})=\bm{\gamma}_{\ell}(\mathbf{c})\odot\mathbf{h}_{\ell}+\bm{\beta}_{\ell}(\mathbf{c})$ [Perez et al., 2018]. This ensures that hidden features and thus the noise predictor $\bm{\epsilon}_{\phi}(\bm{\theta}_{t},\mathbf{y},t)$ remain observation-aware and time-aware at every step, avoiding the signal dilution of naive concatenation.
Given this architecture, to perform amortized inference we draw $N$ initial states $\bm{\theta}_{T}^{(i)}\sim\mathcal{N}(\mathbf{0},\mathbf{I})$ and evolve them backward for $t=T,\ldots,1$ :
$\bm{\theta}_{t-1}^{(i)}=\frac{1}{\sqrt{\alpha_{t}}}\!\left(\bm{\theta}_{t}^{(i)}-\frac{1-\alpha_{t}}{\sqrt{1-\bar{\alpha}_{t}}}\,\hat{\bm{\epsilon}}_{\phi}(\bm{\theta}_{t}^{(i)},\mathbf{y},t)\right)+\sigma_{t}\,\mathbf{z}_{t},$ (16)
where $\mathbf{z}_{t}\sim\mathcal{N}(\mathbf{0},\mathbf{I})$ and $\sigma_{t}=\sqrt{\beta_{t}}$. To balance fidelity to $\mathbf{y}$ with sample diversity, we apply classifier-free guidance by interpolating conditional and unconditional predictions, $\hat{\bm{\epsilon}}_{\phi}=(1+\lambda)\,\bm{\epsilon}_{\phi}(\bm{\theta}_{t},\mathbf{y},t)-\lambda\,\bm{\epsilon}_{\phi}(\bm{\theta}_{t},t),$ with guidance scale $\lambda\geq 0$. The final set $\{\bm{\theta}_{0}^{(i)}\}_{i=1}^{N}$ provides approximate samples from $p(\bm{\theta}\mid\mathbf{y})$ [Ho and Salimans, 2021].